Quadratic Equation
9. The area of a triangle is 20cm^2, and the altitude is 4cm greater than the base. Find the length of the base, to the nearest millimetre.
Thanks!!!

Answer by edjones(8007)

(1/2)ab=A
a=b+4
(1/2)b(b+4)=20
b^2+4b=40
b^2+4b-40=0
b^2+4b =40
b^2+4b+4=40+4 completing the square.
(b+2)^2=44
b+2=+-sqrt(44)
b=-2+-sqrt(11*4)
b=-2+-2sqrt(11)
b=-2+2sqrt(11) cm, b=-2-2sqrt(11) cm
.
Ed